# Framing global structural identifiability in terms of parameter symmetries

### v1

@@ 3.6 Analysing structural identifiability of an epidemiological model
In total, the system has nine rate parameters so that
@@
We implement the CaLinInv-recipe (Algorithm 2) in order to analyse global structural identifiability. Importantly, we compare the outcomes of these calculations to those obtained through the standard differential algebra approach. Also, we generate and visualise parameter symmetries of this model. The underlying calculations were conducted using the open-source symbolic solver SymPy [20]. Details and relevant scripts are available at the public github-repository associated with this project; https://github.com/JohannesBorgqvist/symmetries_and_structural_identifiability.
@@
3.6.1 Finding generating vector fields by solving the linearised symmetry conditions
Starting with the first step in the CaLinInv-recipe, the system of output ODEs is given by
$\dot{y_{E}}=$ $-\frac{\delta y_{E}}{\upsilon}-\frac{k_{E}\mu_{I}y_{I}}{k_{I}}-\frac{k_{E}\dot{y_{I}}}{k_{I}}-\mu_{E}y_{E}+\frac{k_{E}\mu_{I}y_{I}}{\upsilon k_{I}}+\frac{k_{E}\dot{y_{I}}}{\upsilon k_{I}},$ (67)
$\ddot{y_{I}}=$ ${\color[rgb]{1,1,1}\definecolor[named]{pgfstrokecolor}{rgb}{1,1,1}\pgfsys@color@gray@stroke{1}\pgfsys@color@gray@fill{1}-}\beta c\upsilon y_{I}+\frac{\beta\delta y_{E}y_{I}}{k_{E}}-\frac{\beta\mu_{I}y_{I}^{2}}{k_{I}}-\frac{\beta y_{I}\dot{y_{I}}}{k_{I}}+\frac{\delta k_{I}\mu_{S}y_{E}}{k_{E}}-\frac{\delta k_{I}y_{E}\dot{y_{I}}}{k_{E}y_{I}}+\frac{\delta k_{I}\dot{y_{E}}}{k_{E}}$
$-\mu_{I}\mu_{S}y_{I}-\mu_{S}\dot{y_{I}}+\frac{\left(\dot{y_{I}}\right)^{2}}{y_{I}}.$ (68)
$X_{\boldsymbol{\theta}}=\chi_{c}\partial_{c}+\chi_{\beta}\partial_{\beta}+\chi_{\delta}\partial_{\delta}+\chi_{\upsilon}\partial_{\upsilon}+\chi_{\mu_{S}}\partial_{\mu_{S}}+\chi_{\mu_{E}}\partial_{\mu_{E}}+\chi_{\mu_{I}}\partial_{\mu_{I}}+\chi_{k_{E}}\partial_{k_{E}}+\chi_{k_{I}}\partial_{k_{I}},$ (69)
where all infinitesimals are functions of the rate parameters $\boldsymbol{\theta}$ in Eq. (66). The two linearised symmetry conditions defining the generating vector fields can be simplified to the following two equations
$0=$ ${\color[rgb]{1,1,1}\definecolor[named]{pgfstrokecolor}{rgb}{1,1,1}\pgfsys@color@gray@stroke{1}\pgfsys@color@gray@fill{1}-}\beta ck_{E}^{2}k_{I}^{2}\chi_{\upsilon}y_{I}^{2}-\beta\delta k_{I}^{2}\chi_{k_{E}}y_{E}y_{I}^{2}+\beta\upsilon k_{E}^{2}k_{I}^{2}\chi_{c}y_{I}^{2}-\beta k_{E}^{2}k_{I}\chi_{\mu_{I}}y_{I}^{3}+\beta k_{E}^{2}\mu_{I}\chi_{k_{I}}y_{I}^{3}$
$+\beta k_{E}^{2}\chi_{k_{I}}y_{I}^{2}\dot{y_{I}}+\beta k_{E}k_{I}^{2}\chi_{\delta}y_{E}y_{I}^{2}+c\upsilon k_{E}^{2}k_{I}^{2}\chi_{\beta}y_{I}^{2}+\delta k_{E}k_{I}^{3}\chi_{\mu_{S}}y_{E}y_{I}$
$+\delta k_{E}k_{I}^{2}\mu_{S}\chi_{k_{I}}y_{E}y_{I}+\delta k_{E}k_{I}^{2}\chi_{\beta}y_{E}y_{I}^{2}-\delta k_{E}k_{I}^{2}\chi_{k_{I}}y_{E}\dot{y_{I}}+\delta k_{E}k_{I}^{2}\chi_{k_{I}}y_{I}\dot{y_{E}}$
$-\delta k_{I}^{3}\mu_{S}\chi_{k_{E}}y_{E}y_{I}+\delta k_{I}^{3}\chi_{k_{E}}y_{E}\dot{y_{I}}-\delta k_{I}^{3}\chi_{k_{E}}y_{I}\dot{y_{E}}-k_{E}^{2}k_{I}^{2}\mu_{I}\chi_{\mu_{S}}y_{I}^{2}$
$-k_{E}^{2}k_{I}^{2}\mu_{S}\chi_{\mu_{I}}y_{I}^{2}-k_{E}^{2}k_{I}^{2}\chi_{\mu_{S}}y_{I}\dot{y_{I}}-k_{E}^{2}k_{I}\mu_{I}\chi_{\beta}y_{I}^{3}-k_{E}^{2}k_{I}\chi_{\beta}y_{I}^{2}\dot{y_{I}}$
$+k_{E}k_{I}^{3}\mu_{S}\chi_{\delta}y_{E}y_{I}-k_{E}k_{I}^{3}\chi_{\delta}y_{E}\dot{y_{I}}+k_{E}k_{I}^{3}\chi_{\delta}y_{I}\dot{y_{E}},$ (70)
$0=$ ${\color[rgb]{1,1,1}\definecolor[named]{pgfstrokecolor}{rgb}{1,1,1}\pgfsys@color@gray@stroke{1}\pgfsys@color@gray@fill{1}-}\delta k_{I}^{2}\chi_{\upsilon}y_{E}-\upsilon^{2}k_{E}k_{I}\chi_{\mu_{I}}y_{I}+\upsilon^{2}k_{E}\mu_{I}\chi_{k_{I}}y_{I}+\upsilon^{2}k_{E}\chi_{k_{I}}\dot{y_{I}}-\upsilon^{2}k_{I}^{2}\chi_{\mu_{E}}y_{E}$
$-\upsilon^{2}k_{I}\mu_{I}\chi_{k_{E}}y_{I}-\upsilon^{2}k_{I}\chi_{k_{E}}\dot{y_{I}}+\upsilon k_{E}k_{I}\chi_{\mu_{I}}y_{I}-\upsilon k_{E}\mu_{I}\chi_{k_{I}}y_{I}$
$-\upsilon k_{E}\chi_{k_{I}}\dot{y_{I}}-\upsilon k_{I}^{2}\chi_{\delta}y_{E}+\upsilon k_{I}\mu_{I}\chi_{k_{E}}y_{I}+\upsilon k_{I}\chi_{k_{E}}\dot{y_{I}}-k_{E}k_{I}\mu_{I}\chi_{\upsilon}y_{I}$
$-k_{E}k_{I}\chi_{\upsilon}\dot{y_{I}}.$ (71)
The first linearised symmetry condition in Eq. (70) decomposes into subequations based on the products between the various powers of the states and their derivatives according to
$y_{E}\dot{y_{I}}\quad:\quad{0}$ $=-\delta k_{E}k_{I}^{2}\chi_{k_{I}}+\delta k_{I}^{3}\chi_{k_{E}}-k_{E}k_{I}^{3}\chi_{\delta},$ (72)
$y_{I}\dot{y_{E}}\quad:\quad{0}$ $=\delta k_{E}k_{I}^{2}\chi_{k_{I}}-\delta k_{I}^{3}\chi_{k_{E}}+k_{E}k_{I}^{3}\chi_{\delta},$ (73)
$y_{I}\dot{y_{I}}\quad:\quad{0}$ $=k_{E}^{2}k_{I}^{2}\chi_{\mu_{S}},$ (74)
$y_{E}y_{I}\quad:\quad{0}$ $=\delta k_{E}k_{I}^{3}\chi_{\mu_{S}}+\delta k_{E}k_{I}^{2}\mu_{S}\chi_{k_{I}}-\delta k_{I}^{3}\mu_{S}\chi_{k_{E}}+k_{E}k_{I}^{3}\mu_{S}\chi_{\delta},$ (75)
$y_{I}^{2}\quad:\quad{0}$ $=\beta ck_{E}^{2}k_{I}^{2}\chi_{\upsilon}+\beta\upsilon k_{E}^{2}k_{I}^{2}\chi_{c}+c\upsilon k_{E}^{2}k_{I}^{2}\chi_{\beta}-k_{E}^{2}k_{I}^{2}\mu_{I}\chi_{\mu_{S}}-k_{E}^{2}k_{I}^{2}\mu_{S}\chi_{\mu_{I}},$ (76)
$y_{I}^{2}\dot{y_{I}}\quad:\quad{0}$ $=\beta k_{E}^{2}\chi_{k_{I}}-k_{E}^{2}k_{I}\chi_{\beta},$ (77)
$y_{E}y_{I}^{2}\quad:\quad{0}$ $=-\beta\delta k_{I}^{2}\chi_{k_{E}}+\beta k_{E}k_{I}^{2}\chi_{\delta}+\delta k_{E}k_{I}^{2}\chi_{\beta},$ (78)
$y_{I}^{3}\quad:\quad{0}$ $=-\beta k_{E}^{2}k_{I}\chi_{\mu_{I}}+\beta k_{E}^{2}\mu_{I}\chi_{k_{I}}-k_{E}^{2}k_{I}\mu_{I}\chi_{\beta}.$ (79)
Similarly, the second linearised symmetry condition in Eq. (71) decomposes into
$\dot{y_{I}}\quad:\quad{0}$ $=\upsilon^{2}k_{E}\chi_{k_{I}}-\upsilon^{2}k_{I}\chi_{k_{E}}-\upsilon k_{E}\chi_{k_{I}}+\upsilon k_{I}\chi_{k_{E}}-k_{E}k_{I}\chi_{\upsilon},\quad\quad\;\;$ (80)
$y_{E}\quad:\quad{0}$ $=\delta k_{I}^{2}\chi_{\upsilon}-\upsilon^{2}k_{I}^{2}\chi_{\mu_{E}}-\upsilon k_{I}^{2}\chi_{\delta},$ (81)
$\quad+\upsilon k_{E}k_{I}\chi_{\mu_{I}}-\upsilon k_{E}\mu_{I}\chi_{k_{I}}+\upsilon k_{I}\mu_{I}\chi_{k_{E}}-k_{E}k_{I}\mu_{I}\chi_{\upsilon}.$ (82)
${\tiny M=\begin{pmatrix}0&0&-k_{E}k_{I}^{3}&0&0&0&0&-\delta k_{E}k_{I}^{2}&\delta k_{I}^{3}\\ 0&0&k_{E}k_{I}^{3}&0&0&0&0&\delta k_{E}k_{I}^{2}&-\delta k_{I}^{3}\\ 0&0&0&0&-k_{E}^{2}k_{I}^{2}&0&0&0&0\\ 0&0&k_{E}k_{I}^{3}\mu_{S}&0&\delta k_{E}k_{I}^{3}&0&0&\delta k_{E}k_{I}^{2}\mu_{S}&-\delta k_{I}^{3}\mu_{S}\\ \beta\upsilon k_{E}^{2}k_{I}^{2}&c\upsilon k_{E}^{2}k_{I}^{2}&0&\beta ck_{E}^{2}k_{I}^{2}&-k_{E}^{2}k_{I}^{2}\mu_{I}&0&-k_{E}^{2}k_{I}^{2}\mu_{S}&0&0\\ 0&-k_{E}^{2}k_{I}&0&0&0&0&0&\beta k_{E}^{2}&0\\ 0&\delta k_{E}k_{I}^{2}&\beta k_{E}k_{I}^{2}&0&0&0&0&0&-\beta\delta k_{I}^{2}\\ 0&-k_{E}^{2}k_{I}\mu_{I}&0&0&0&0&-\beta k_{E}^{2}k_{I}&\beta k_{E}^{2}\mu_{I}&0\\ 0&0&0&-k_{E}k_{I}&0&0&0&\upsilon^{2}k_{E}-\upsilon k_{E}&-\upsilon^{2}k_{I}+\upsilon k_{I}\\ 0&0&-\upsilon k_{I}^{2}&\delta k_{I}^{2}&0&-\upsilon^{2}k_{I}^{2}&0&0&0\\ 0&0&0&-k_{E}k_{I}\mu_{I}&0&0&-\upsilon^{2}k_{E}k_{I}+\upsilon k_{E}k_{I}&\upsilon^{2}k_{E}\mu_{I}-\upsilon k_{E}\mu_{I}&-\upsilon^{2}k_{I}\mu_{I}+\upsilon k_{I}\mu_{I}\end{pmatrix}.}$ (83)
$\mathcal{N}(M)\coloneqq\left\{\boldsymbol{\chi}=\begin{pmatrix}\chi_{c}\\ \chi_{\beta}\\ \chi_{\delta}\\ \chi_{\upsilon}\\ \chi_{\mu_{S}}\\ \chi_{\mu_{E}}\\ \chi_{\mu_{I}}\\ \chi_{k_{I}}\\ \chi_{k_{E}}\end{pmatrix}\in\mathbb{R}^{9}:\boldsymbol{\chi}\in\mathrm{Span}\left\{\begin{pmatrix}-c\upsilon\\ \beta\\ -\delta\\ \upsilon\left(\upsilon-1\right)\\ 0\\ \delta\\ 0\\ k_{I}\\ 0\end{pmatrix},\begin{pmatrix}c\left(\upsilon-1\right)\\ 0\\ \delta\\ \upsilon\left(1-\upsilon\right)\\ 0\\ -\delta\\ 0\\ 0\\ k_{E}\end{pmatrix}\right\}\right\}.$ (84)
As such, we consider the following parameter infinitesimals
$\begin{pmatrix}\chi_{c}\\ \chi_{\beta}\\ \chi_{\delta}\\ \chi_{\upsilon}\\ \chi_{\mu_{S}}\\ \chi_{\mu_{E}}\\ \chi_{\mu_{I}}\\ \chi_{k_{I}}\\ \chi_{k_{E}}\end{pmatrix}=\alpha_{1}\begin{pmatrix}-c\upsilon\\ \beta\\ -\delta\\ \upsilon\left(\upsilon-1\right)\\ 0\\ \delta\\ 0\\ k_{I}\\ 0\end{pmatrix}+\alpha_{2}\begin{pmatrix}c\left(\upsilon-1\right)\\ 0\\ \delta\\ \upsilon\left(1-\upsilon\right)\\ 0\\ -\delta\\ 0\\ 0\\ k_{E}\end{pmatrix}=\begin{pmatrix}(\alpha_{2}-\alpha_{1})c\upsilon-c\alpha_{2}\\ \alpha_{1}\beta\\ (\alpha_{2}-\alpha_{1})\delta\\ (\alpha_{1}-\alpha_{2})\upsilon\left(\upsilon-1\right)\\ 0\\ -(\alpha_{2}-\alpha_{1})\delta\\ 0\\ \alpha_{1}k_{I}\\ \alpha_{2}k_{E}\end{pmatrix},$ (85)
$\begin{split}X_{\boldsymbol{\theta}}=&\,\,[(\alpha_{2}-\alpha_{1})c\upsilon-\alpha_{2}c]\partial_{c}+\alpha_{1}\beta\partial_{\beta}+(\alpha_{2}-\alpha_{1})\delta\partial_{\delta}+(\alpha_{1}-\alpha_{2})\upsilon\left(\upsilon-1\right)\partial_{\upsilon}\\ &\,\,\,\,\,-(\alpha_{2}-\alpha_{1})\delta\partial_{\mu_{E}}+\alpha_{1}k_{I}\partial_{k_{I}}+\alpha_{2}k_{E}\partial_{k_{E}}.\end{split}$ (86)
3.6.2 Elucidating structural identifiability by calculating the universal parameter invariants
The structurally identifiable quantities are given by universal parameter invariants. Thus, we need to find parameter invariants that are independent of the arbitrary coefficients $\alpha_{1}$ and $\alpha_{2}$ appearing in Eq. (86). To this end, let $s$ be a parameter that parametrises the parameter invariants $I$ of interest according to
$I(s)=I(c(s),\beta(s),\delta(s),\upsilon(s),\mu_{S}(s),\mu_{E}(s),\mu_{I}(s),k_{E}(s),k_{I}(s)).$ (87)
Then, the characteristic equations are given by
$\frac{\mathrm{d}c}{\mathrm{d}s}$ $=\chi_{c}=(\alpha_{2}-\alpha_{1})c\upsilon-\alpha_{2}c,$ (88)
$\frac{\mathrm{d}\beta}{\mathrm{d}s}$ $=\chi_{b}=\alpha_{1}\beta,$ (89)
$\frac{\mathrm{d}\delta}{\mathrm{d}s}$ $=\chi_{\delta}=(\alpha_{2}-\alpha_{1})\delta,$ (90)
$\frac{\mathrm{d}\upsilon}{\mathrm{d}s}$ $=\chi_{\upsilon}=(\alpha_{1}-\alpha_{2})\upsilon\left(\upsilon-1\right),$ (91)
$\frac{\mathrm{d}\mu_{S}}{\mathrm{d}s}$ $=\chi_{\mu_{S}}=0,$ (92)
$\frac{\mathrm{d}\mu_{E}}{\mathrm{d}s}$ $=\chi_{\mu_{E}}=-(\alpha_{2}-\alpha_{1})\delta,$ (93)
$\frac{\mathrm{d}\mu_{I}}{\mathrm{d}s}$ $=\chi_{\mu_{I}}=0,$ (94)
$\frac{\mathrm{d}k_{E}}{\mathrm{d}s}$ $=\chi_{k_{E}}=\alpha_{2}k_{E},$ (95)
$\frac{\mathrm{d}k_{I}}{\mathrm{d}s}$ $=\chi_{k_{I}}=\alpha_{1}k_{I}.$ (96)
The universal parameter invariants are first integrals of Eqs. (88)-(96) that are independent of the arbitrary coefficients $\alpha_{1}$ and $\alpha_{2}$. By Eqs. (92) and (94), two universal parameter invariants are given by
$I_{1}=\mu_{I},\quad I_{2}=\mu_{S},$ (97)
All of the remaining rate parameters are unidentifiable, and next we set out to find the remaining universal parameter invariants. Combining Eqs. (90) and (93), we obtain
$\frac{\mathrm{d}\delta}{\mathrm{d}\mu_{E}}=-1,$ (98)
and the corresponding universal parameter invariant is given by
$I_{3}=\delta+\mu_{E}.$ (99)
Combining Eqs. (89) and (96), we obtain
$\frac{\mathrm{d}\beta}{\mathrm{d}k_{I}}=\frac{\beta}{k_{I}},$ (100)
and the corresponding universal parameter invariant is given by
$I_{4}=\frac{\beta}{k_{I}}.$ (101)
Combining Eqs. (90) and (91), we obtain
$\frac{\mathrm{d}\upsilon}{\mathrm{d}\delta}=\frac{\upsilon(1-\upsilon)}{\delta},$ (102)
and the corresponding universal parameter invariant is given by
$I_{5}=\frac{\upsilon}{\delta(1-\upsilon)}.$ (103)
These five invariants are simple to calculate as it is obvious how the arbitrary coefficients $\alpha_{1}$ and $\alpha_{2}$ are eliminated. For the two remaining invariants involving the parameters $c$ and $k_{E}$ some algebraic manipulations are required to eliminate the coefficients $\alpha_{1}$ and $\alpha_{2}$ in order to find the corresponding universal parameter invariants. Starting with the parameter $c$, we consider the product $c\upsilon$. Using the characteristic equations in Eqs. (88) and (91) yields
$\frac{\mathrm{d}(c\upsilon)}{\mathrm{d}s}=\upsilon\chi_{c}+c\chi_{\upsilon}=-\alpha_{1}(c\upsilon),$ (104)
and combining the resulting characteristic equation for $c\upsilon$ with that in Eq. (89) yields
$\frac{\mathrm{d}(c\upsilon)}{\mathrm{d}\beta}=-\frac{(c\upsilon)}{\beta}.$ (105)
$I_{6}=\beta{c}\upsilon.$ (106)
Finally, for the parameter $k_{E}$ we consider the quotient $\delta/k_{E}$. Using the characteristic equations in Eqs. (90) and (95) we obtain the following characteristic equation for the quotient $\delta/k_{E}$
$\frac{\mathrm{d}}{\mathrm{d}s}\left(\frac{\delta}{k_{E}}\right)=\frac{k_{E}\chi_{\delta}-\delta\chi_{k_{E}}}{k_{E}^{2}}=-\alpha_{1}\left(\frac{\delta}{k_{E}}\right),$ (107)
and combining this characteristic equation with that in Eq. (89) yields
$\frac{\mathrm{d}(\delta/k_{E})}{\mathrm{d}\beta}=-\frac{(\delta/k_{E})}{\beta}.$ (108)
The last universal parameter invariant is therefore
$I_{7}=\frac{\beta\delta}{k_{E}}.$ (109)
The seven universal parameter invariants $I_{1}$, $I_{2}$, $I_{3}$, $I_{4}$, $I_{5}$, $I_{6}$ and $I_{7}$ correspond to the same parameter manifold found by Renardy et al. [16] using the standard differential algebra approach.
3.6.3 Generating and visualising parameter symmetries of the SEI model
We generate parameter symmetries of the SEI model using the vector field $X_{\boldsymbol{\theta}}$ in Eq. (86). Specifically, these symmetries $\Gamma_{\varepsilon}^{\boldsymbol{\theta}}$ generate transformed parameter vectors according to $\Gamma_{\varepsilon}^{\boldsymbol{\theta}}:\boldsymbol{\theta}\mapsto\hat{\boldsymbol{\theta}}(\varepsilon)$ that are given by
@@ 3.6.3 Generating and visualising parameter symmetries of the SEI model
$\frac{\mathrm{d}\hat{c}}{\mathrm{d}\varepsilon}$ $=(\alpha_{2}-\alpha_{1})\hat{c}\hat{\upsilon}-\alpha_{2}\hat{c},\quad$ $\hat{c}(\varepsilon=0)=c,$ (111)
$\frac{\mathrm{d}\hat{\beta}}{\mathrm{d}\varepsilon}$ $=\alpha_{1}\hat{\beta},\quad$ $\hat{\beta}(\varepsilon=0)=\beta,$ (112)
$\frac{\mathrm{d}\hat{\delta}}{\mathrm{d}\varepsilon}$ $=(\alpha_{2}-\alpha_{1})\hat{\delta},\quad$ $\hat{\delta}(\varepsilon=0)=\delta,$ (113)
@@
$\frac{\mathrm{d}\hat{\mu}_{E}}{\mathrm{d}\varepsilon}$ $=-(\alpha_{2}-\alpha_{1})\hat{\delta},\quad$ $\hat{\mu}_{E}(\varepsilon=0)=\mu_{E},$ (116)
@@
This system of ODEs is readily solved numerically in order to characterise the action of any specific symmetry defined by specific choices of the coefficients $\alpha_{1}$ and $\alpha_{2}$. Since the parameter space of the SEI model is nine-dimensional, we visualise the action of the parameter symmetry of interest in two- and three-dimensional subspaces. Specifically, we have numerically generated six parameter vectors as starting points illustrated by diamonds in order to plot the corresponding transformed parameter vectors $\hat{\boldsymbol{\theta}}(\varepsilon)$ in Eq. (110).
@@
Fig. 1: Two-dimensional projections of the action of a parameter symmetry $\Gamma_{\varepsilon}^{\boldsymbol{\theta}}$ of the SEI model. The action of the symmetry $\Gamma_{\varepsilon}^{\boldsymbol{\theta}}$ generated by solving the system of ODEs in Eqs. (111)-(119) for the coefficients $(\alpha_{1},\alpha_{2})=(2,1)$ on six parameters illustrated by diamonds is visualised in three different two-dimensional subspaces of the nine-dimensional parameter space of the SEI model. These subspaces are (A) $(\mu_{E},\delta)$ for which the invariant $I_{3}=\delta+\mu_{E}$ is preserved, (B) $(k_{I},\beta)$ for which the invariant $I_{4}=\beta/k_{I}$ is preserved and (C) $(\delta,\upsilon)$ for which the invariant $I_{5}=\upsilon/(\delta(1-\upsilon))$ is preserved.

### v3

@@ 4.5 Analysing local structural identifiability of an epidemiological model
In total, the system has nine parameters so that
@@
The system of output ODEs is given by
$\dfrac{\mathrm{d}y_{E}}{\mathrm{d}t}=$ $-\frac{\delta y_{E}}{\upsilon}-\frac{k_{E}\mu_{I}y_{I}}{k_{I}}-\frac{k_{E}}{k_{I}}\dfrac{\mathrm{d}y_{I}}{\mathrm{d}t}-\mu_{E}y_{E}+\frac{k_{E}\mu_{I}y_{I}}{\upsilon k_{I}}+\frac{k_{E}}{\upsilon k_{I}}\dfrac{\mathrm{d}y_{I}}{\mathrm{d}t},$ (63)
$\dfrac{\mathrm{d}^{2}y_{I}}{\mathrm{d}t^{2}}=$ ${\color[rgb]{1,1,1}-}\beta c\upsilon y_{I}+\frac{\beta\delta y_{E}y_{I}}{k_{E}}-\frac{\beta\mu_{I}y_{I}^{2}}{k_{I}}-\frac{\beta y_{I}}{k_{I}}\dfrac{\mathrm{d}y_{I}}{\mathrm{d}t}+\frac{\delta k_{I}\mu_{S}y_{E}}{k_{E}}-\frac{\delta k_{I}y_{E}}{k_{E}y_{I}}\dfrac{\mathrm{d}y_{I}}{\mathrm{d}t}+\frac{\delta k_{I}}{k_{E}}\dfrac{\mathrm{d}y_{E}}{\mathrm{d}t}$
$-\mu_{I}\mu_{S}y_{I}-\mu_{S}\dfrac{\mathrm{d}y_{I}}{\mathrm{d}t}+\frac{1}{y_{I}}\left(\dfrac{\mathrm{d}y_{I}}{\mathrm{d}t}\right)^{2}.$ (64)
Step 2: Linearised symmetry conditions
The two linearised symmetry conditions can be simplified to the following two equations
$0=$ ${\color[rgb]{1,1,1}-}\beta ck_{E}^{2}k_{I}^{2}\chi_{\upsilon}y_{I}^{2}-\beta\delta k_{I}^{2}\chi_{k_{E}}y_{E}y_{I}^{2}+\beta\upsilon k_{E}^{2}k_{I}^{2}\chi_{c}y_{I}^{2}-\beta k_{E}^{2}k_{I}\chi_{\mu_{I}}y_{I}^{3}+\beta k_{E}^{2}\mu_{I}\chi_{k_{I}}y_{I}^{3}$
$+\beta k_{E}^{2}\chi_{k_{I}}y_{I}^{2}\dfrac{\mathrm{d}y_{I}}{\mathrm{d}t}+\beta k_{E}k_{I}^{2}\chi_{\delta}y_{E}y_{I}^{2}+c\upsilon k_{E}^{2}k_{I}^{2}\chi_{\beta}y_{I}^{2}+\delta k_{E}k_{I}^{3}\chi_{\mu_{S}}y_{E}y_{I}$
$+\delta k_{E}k_{I}^{2}\mu_{S}\chi_{k_{I}}y_{E}y_{I}+\delta k_{E}k_{I}^{2}\chi_{\beta}y_{E}y_{I}^{2}-\delta k_{E}k_{I}^{2}\chi_{k_{I}}y_{E}\dfrac{\mathrm{d}y_{I}}{\mathrm{d}t}+\delta k_{E}k_{I}^{2}\chi_{k_{I}}y_{I}\dfrac{\mathrm{d}y_{E}}{\mathrm{d}t}$
$-\delta k_{I}^{3}\mu_{S}\chi_{k_{E}}y_{E}y_{I}+\delta k_{I}^{3}\chi_{k_{E}}y_{E}\dfrac{\mathrm{d}y_{I}}{\mathrm{d}t}-\delta k_{I}^{3}\chi_{k_{E}}y_{I}\dfrac{\mathrm{d}y_{E}}{\mathrm{d}t}-k_{E}^{2}k_{I}^{2}\mu_{I}\chi_{\mu_{S}}y_{I}^{2}$
$-k_{E}^{2}k_{I}^{2}\mu_{S}\chi_{\mu_{I}}y_{I}^{2}-k_{E}^{2}k_{I}^{2}\chi_{\mu_{S}}y_{I}\dfrac{\mathrm{d}y_{I}}{\mathrm{d}t}-k_{E}^{2}k_{I}\mu_{I}\chi_{\beta}y_{I}^{3}-k_{E}^{2}k_{I}\chi_{\beta}y_{I}^{2}\dfrac{\mathrm{d}y_{I}}{\mathrm{d}t}$
$+k_{E}k_{I}^{3}\mu_{S}\chi_{\delta}y_{E}y_{I}-k_{E}k_{I}^{3}\chi_{\delta}y_{E}\dfrac{\mathrm{d}y_{I}}{\mathrm{d}t}+k_{E}k_{I}^{3}\chi_{\delta}y_{I}\dfrac{\mathrm{d}y_{E}}{\mathrm{d}t},$ (65)
$0=$ ${\color[rgb]{1,1,1}-}\delta k_{I}^{2}\chi_{\upsilon}y_{E}-\upsilon^{2}k_{E}k_{I}\chi_{\mu_{I}}y_{I}+\upsilon^{2}k_{E}\mu_{I}\chi_{k_{I}}y_{I}+\upsilon^{2}k_{E}\chi_{k_{I}}\dfrac{\mathrm{d}y_{I}}{\mathrm{d}t}-\upsilon^{2}k_{I}^{2}\chi_{\mu_{E}}y_{E}$
$-\upsilon^{2}k_{I}\mu_{I}\chi_{k_{E}}y_{I}-\upsilon^{2}k_{I}\chi_{k_{E}}\dfrac{\mathrm{d}y_{I}}{\mathrm{d}t}+\upsilon k_{E}k_{I}\chi_{\mu_{I}}y_{I}-\upsilon k_{E}\mu_{I}\chi_{k_{I}}y_{I}$
$-\upsilon k_{E}\chi_{k_{I}}\dfrac{\mathrm{d}y_{I}}{\mathrm{d}t}-\upsilon k_{I}^{2}\chi_{\delta}y_{E}+\upsilon k_{I}\mu_{I}\chi_{k_{E}}y_{I}+\upsilon k_{I}\chi_{k_{E}}\dfrac{\mathrm{d}y_{I}}{\mathrm{d}t}-k_{E}k_{I}\mu_{I}\chi_{\upsilon}y_{I}$
$-k_{E}k_{I}\chi_{\upsilon}\dfrac{\mathrm{d}y_{I}}{\mathrm{d}t}.$ (66)
@@
The family of generating vector fields solving these linearised symmetry conditions is given by
[…]
where $\alpha_{1}$ and $\alpha_{2}$ are two arbitrary coefficients. Using this family of vector fields, we generate parameter symmetries of the SEI model. Specifically, these symmetries $\Gamma_{\varepsilon}^{\boldsymbol{\theta}}$ generate transformed parameter vectors according to $\Gamma_{\varepsilon}^{\boldsymbol{\theta}}:\boldsymbol{\theta}\mapsto\hat{\boldsymbol{\theta}}(\varepsilon)$ that are given by
@@
$\frac{\mathrm{d}\hat{c}}{\mathrm{d}\varepsilon}$ $=(\alpha_{2}-\alpha_{1})\hat{c}\hat{\upsilon}-\alpha_{2}\hat{c}$$\quad$ $\hat{c}(\varepsilon=0)=c,$ (69)
$\frac{\mathrm{d}\hat{\beta}}{\mathrm{d}\varepsilon}$ $=\alpha_{1}\hat{\beta}$$\quad$ $\hat{\beta}(\varepsilon=0)=\beta,$ (70)
$\frac{\mathrm{d}\hat{\delta}}{\mathrm{d}\varepsilon}$ $=(\alpha_{2}-\alpha_{1})\hat{\delta}$$\quad$ $\hat{\delta}(\varepsilon=0)=\delta,$ (71)
@@
$\frac{\mathrm{d}\hat{\mu}_{E}}{\mathrm{d}\varepsilon}$ $=-(\alpha_{2}-\alpha_{1})\hat{\delta}$$\quad$ $\hat{\mu}_{E}(\varepsilon=0)=\mu_{E},$ (74)
@@
This system of ODEs is readily solved numerically in order to characterise the action of any specific symmetry defined by specific choices of the coefficients $\alpha_{1}$ and $\alpha_{2}$. Since the parameter space of the SEI model is nine-dimensional, we visualise the action of the parameter symmetry of interest in two- and three-dimensional subspaces. Specifically, we have numerically generated six parameter vectors as starting points illustrated by diamonds in order to plot the corresponding transformed parameter vectors $\hat{\boldsymbol{\theta}}(\varepsilon)$ in Eq. (68), and relevant figures are presented in the next step of the CaLinInv recipe. The interpretation of these curves in parameter space is that the illustrated parameters have indistinguishable model behaviour.
Step 3: Universal Invariants
The following seven universal parameter invariants were found (for details, see the Appendices) using the method of characteristics:
$I_{1}$ $=\mu_{I}\,$, (78)
$I_{2}$ $=\mu_{S}\,$, (79)
$I_{3}$ $=\delta+\mu_{E}\,$, (80)
$I_{4}$ $=\frac{\beta}{k_{I}}\,$, (81)
$I_{5}$ $=\frac{\upsilon}{\delta(1-\upsilon)}\,$, (82)
$I_{6}$ $=\beta{c}\upsilon\,$, (83)
$I_{7}$ $=\frac{\beta\delta}{k_{E}}\,.$ (84)
These universal parameter invariants correspond to the same globally structurally identifiable parameter combinations found by Renardy et al. [16] using the standard differential algebra approach. Since these universal parameter invariants also correspond to the locally structurally identifiable parameter combinations (Cor. 1), the locally and globally structurally identifiable parameter combinations are the same in the case of the SEI model. By solving the ODE system in Eqs. (69)–(77) for the arbitrary coefficients $(\alpha_{1},\alpha_{2})=(2,1)$, we visualise the action of the corresponding parameter symmetry and illustrate that it preserves all seven universal parameter invariants.
@@
Figure 1: Two-dimensional projections of the action of a parameter symmetry $\Gamma_{\varepsilon}^{\boldsymbol{\theta}}$ of the SEI model. The action of the symmetry $\Gamma_{\varepsilon}^{\boldsymbol{\theta}}$ generated by solving the system of ODEs in Eqs. (69)-(77) for the coefficients $(\alpha_{1},\alpha_{2})=(2,1)$ on six parameters illustrated by diamonds is visualised in three different two-dimensional subspaces of the nine-dimensional parameter space of the SEI model. The interpretation of these curves in parameter space is that the illustrated parameters have indistinguishable model behaviour. These subspaces are (A) $(\mu_{E},\delta)$ for which the invariant $I_{3}=\delta+\mu_{E}$ is preserved, (B) $(k_{I},\beta)$ for which the invariant $I_{4}=\beta/k_{I}$ is preserved and (C) $(\delta,\upsilon)$ for which the invariant $I_{5}=\upsilon/(\delta(1-\upsilon))$ is preserved.